# GARCH copulas and GARCH-mimicking copulas

### v1

@@ Abstract
The bivariate copulas that describe the dependencies and partial dependencies of lagged variables in strictly stationary, first-order GARCH-type processes are investigated. It is shown that the copulas of symmetric GARCH processes are jointly symmetric but non-exchangeable, while the copulas of processes with symmetric innovation distributions and asymmetric leverage effects have weaker h-symmetry; copulas with asymmetric innovation distributions have neither form of symmetry. Since the actual copulas are typically inaccessible, due to the unknown functional forms of the marginal distributions of GARCH processes, methods of mimicking them are proposed. These rely on constructions that combine standard bivariate copulas for positive dependence with two uniformity-preserving transformations known as v-transforms. A variety of new copulas are introduced and the ones providing the best fit to simulated data from GARCH-type processes are identified. A method of constructing tractable simplified d-vines using linear v-transforms is described and shown to coincide with the vt-d-vine model when the two v-transforms are identical.
@@ 1 Introduction
The ARCH process (Engle,, 1982), the GARCH process (Bollerslev,, 1986) and a whole lexicon of related GARCH-type models have proved to be remarkably successful models for time series displaying stochastic volatility, such as asset price return series. In this paper we analyse the serial dependence structure of first-order GARCH-type models in terms of the implied bivariate copulas that describe the dependence of pairs of variables at different lags. Our aim is to understand how we could mimic such dependence structures with appropriately chosen bivariate copulas arranged in d-vines; we do this with a view to developing models that might offer additional flexibility in practice. The approach we take recognizes that uniformity-preserving transformations, known as v-transforms (McNeil,, 2021), are embedded in the implied copulas of classical GARCH models and suggests that they could be used as explicit modelling elements to construct vines for applications.
The essential mechanism of a GARCH-type model is simple and intuitive; the conditional distribution from which a new observation is drawn has a standard deviation that is a function of the size of previous returns, so that large market moves have a tendency to beget further large market moves and create periods of volatility. The ARCH and GARCH prototypes of Engle, (1982) and Bollerslev, (1986) have inspired a vast class of related GARCH-type models, including the GJR-GARCH model, which incorporates leverage effects to reflect different contributions to volatility by returns of differing signs (Glosten et al.,, 1993), power-GARCH models (Ding et al.,, 1993) and exponential GARCH models (Nelson,, 1991), to name a but a few of the more important extensions. A number of comprehensive surveys of the extended GARCH family have been published including Bollerslev et al., (1992) Bollerslev et al., (1994) and Shephard, (1996).
@@
Notwithstanding the appealing basic mechanism for explaining volatility, GARCH-type models are also something of a black box. The stationary distribution is not generally known in closed form, although it is known that many models gives rise to distributions with Pareto tails (Mikosch and Stărică,, 2000). In particular, the serial dependence structure is not well understood. If $(X_{t})_{t\in\mathbb{Z}}$ follows a stationary GARCH-type model then the distribution of $(X_{1},\ldots,X_{n})^{\top}$ for $n\geqslant 2$ must have a copula, but this copula also has no simple, accessible closed form.
It may be asked why we would want to mimic GARCH-type models with vine copulas, given that GARCH models have proved themselves in practice over four decades. While many asset return series are well modelled by GARCH, we do encounter series where a power-lawed marginal distribution is sub-optimal for the data; see, for example, Bladt and McNeil, 2022b . For such datasets it would be convenient to be able to construct models that combined bespoke choices of marginal distribution, or non-parametric margins, with a copula that could mimic the kind of serial dependence that is implicit in the GARCH family. If we understand the types of copula that are implicit in the GARCH family, it may also be possible to improve upon these copulas by adding further flexibility.
@@
Vine copulas also have a large literature, including Joe, (1996, 1997), Bedford and Cooke, (2001, 2002), Kurowicka and Cooke, (2006), Aas et al., (2009) and Smith et al., (2010). The most suitable vine structure for a time series is the d-vine, which is able to describe strict stationarity of a random vector under additional translation-invariance restrictions on the vine structure (Nagler et al.,, 2022). Markov copula models (Darsow et al.,, 1992; Chen and Fan,, 2006; Beare,, 2010; Domma et al.,, 2009) are simple examples of stationary first-order d-vines and a number of authors have written on their higher-order extensions (Ibragimov,, 2009; Beare and Seo,, 2015; Brechmann and Czado,, 2015; Loaiza-Maya et al.,, 2018; Nagler et al.,, 2022). The last-named paper introduces the terminology s-vine for stationary multivariate vines, which includes the stationary univariate d-vine as a special case.
@@
The joint density of the distribution of $(X_{1},\ldots,X_{n})^{\top}$ in a strictly stationary process, such as a stationary model of GARCH type, can be decomposed as a d-vine. In other words we may write
$f_{X_{1},\ldots,X_{n}}(x_{1},\ldots,x_{n})=\prod_{i=1}^{n}f_{X}(x_{i})\prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}(F_{X_{1}|X_{2},\ldots,X_{k}}(x_{j-k}\mid x_{j-k+1},\ldots,x_{j-1}),\\ F_{X_{k+1}|X_{2},\ldots,X_{k}}(x_{j}\mid x_{j-k+1},\ldots,x_{j-1})\mid x_{j-k+1},\ldots,x_{j-1})$ (1)
@@
where $f_{X}$ denotes the marginal density, $F_{X_{i}|X_{2},\ldots,X_{k}}$ denotes the conditional distribution function of $X_{i}$ given $X_{2},\ldots,X_{k}$ and $c_{k}(u,v\mid x_{2},\ldots,x_{k})$ is the density of the copula of the conditional distribution of $(X_{1},X_{k+1})^{\top}$ given $X_{2}=x_{2},\ldots,X_{k}=x_{k}$. In the case $k=1$ the conditional distribution is simply the marginal distribution $F_{X}$ and $c_{1}$ is simply the density of the copula of the unconditional distribution of $(X_{1},X_{2})^{\top}$. When the process is first-order Markov, all of the terms involving $c_{k}(u,v\mid\cdot)$ for $k>1$ disappear, but this will only apply to the ARCH(1)-type models among the models we analyse.
@@
Although the densities of GARCH-type models may be written as d-vines, they may not in general be written as so-called simplified d-vines, that is models in which the conditional copula densities $c_{k}$ do not depend explicitly on the values $x_{j-k+1},\ldots,x_{j-1}$ in equation (1) for $k\geqslant 2$. When d-vines are used in practical modelling applications they are almost always simplified d-vines. However, as pointed out by a number of authors (Haff et al.,, 2010; Stöber et al.,, 2013; Spanhel and Kurz,, 2019; Mroz et al.,, 2021), these may not be able to capture all the features present in an arbitrary joint density, such as that of a GARCH-type model. In the course of our analysis we return to this question.
We confine our attention to first-order GARCH-type models. In practice models of order (1,1) often work very well and adding higher-order terms yields only minor improvements. We also look in some detail at ARCH(1)-type models for two reasons: they are first-order Markov models and therefore particularly easy to analyse; they give rise to the essential copula forms that are encountered in the more general class, as will be seen in this paper.
Loaiza-Maya et al., (2018) have previously investigated the copula of the classical ARCH(1) process. Since this model is first-order Markov, its serial dependence may be fully described by a first-order d-vine with a single bivariate copula, namely the copula of the distribution of two successive values $(X_{1},X_{2})^{\top}$. Using a simulation approach, they identify the cross-shaped nature of the copula density and suggest that mixtures of positive dependence and negative dependence copulas (such as a t copula and a 90-degree rotated t copula) can model the dependence structure effectively. Zhao et al., (2022) argue that t copulas in d-vines can approximate the dependence structure of GARCH(1,1) and GJR-GARCH(1,1) models. We will find some evidence in support of t copulas but also argue that they miss some important asymmetries.
@@
We make a number of original contributions. First, we analyse the structure of the bivariate copula densities $c_{k}$ in (1) for classical GARCH-type models; we show that symmetric models give rise to jointly symmetric copulas while models with symmetric innovation distributions and asymmetric volatility have copulas with weaker symmetry, which we call h-symmetry. We confirm the theoretical insights by using numerical approximations to provide accurate contour plots of the copula densities. These reveal hitherto overlooked features, such as the non-exchangeability of the copulas. We also examine the effects of leverage, asymmetric innovations and autoregressive conditional mean terms on the copulas and on their symmetries.
@@
Second, we show how uniformity-preserving transformations known as v-transforms (McNeil,, 2021) play an implicit role in the copulas of classic symmetric models and in models with certain forms of asymmetry. We investigate methods of constructing copulas with v-transforms to mimic the behaviour of the copulas implied by GARCH-type models. Using simulated data we identify the copulas based on v-transforms that provide the best fit to simulated data and show that these typically do better than mixture copulas.
Finally, we draw conclusions with respect to the construction of d-vine copulas to mimic the serial-dependence structure of GARCH-type models. We also show how d-vine models with appropriately chosen copulas can be related to the vt-ARMA and vt-d-vine models proposed by McNeil, (2021) and Bladt and McNeil, 2022b .
2 Preliminary results on symmetry
2.1 Definitions
Symmetry plays an important role in this paper and we give a definition to make our use of various symmetry concepts clear; parts (iii) and (iv) are from Nelsen, (2006).
Definition 1.
Let $(Y,Z)^{\top}$ be a generic random vector taking values in $\mathbb{R}^{2}$.
  1. (i)
    If $(Y,Z-b)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(Y,b-Z)^{\top}$ for some $b\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is h-symmetric about $b$ (symmetric about a horizontal axis at $b$).
  2. (ii)
    If $(Y-a,Z)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(a-Y,Z)^{\top}$ for some $a\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is v-symmetric about $a$ (symmetric about a vertical axis at $a$).
  3. (iii)
    If $(Y-a,Z-b)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(a-Y,b-Z)^{\top}$ for $a,b\in\mathbb{R}$ we say that $(Y,Z)^{\top}$ is radially symmetry about $(a,b)$.
  4. (iv)
    If $(Y,Z)^{\top}$ has all three of the above forms of symmetry then we say that it is jointly symmetric about $(a,b)$.
Remarks 1.
  1. 1.
    When $(Y,Z)^{\top}$ is h-symmetric about $b$, the marginal distribution of $Z$ is symmetric about $b$; when $(Y,Z)^{\top}$ is v-symmetric about $a$, the marginal distribution of $Y$ is symmetric about $a$. The reverse implications are not true.
  2. 2.
    It may be noted that we can always shift the distribution of $(Y,Z)^{\top}$ in two-dimensional space so that axes or centres of symmetry are at zero or the origin. Where $a$ and $b$ are not relevant to the result, we drop the explicit reference to them and just refer to h-symmetry, v-symmetry, radial symmetry and joint symmetry.
  3. 3.
    H-symmetry and v-symmetry together imply radial symmetry and therefore joint symmetry. However radial symmetry does not necessarily imply h-symmetry and v-symmetry, although it does imply symmetry of both marginal distributions.
The symmetries in Definition 1 have the following implications for the copula of $(Y,Z)^{\top}$.
Lemma 1.
  1. (i)
    If $(Y,Z)^{\top}$ is h-symmetric then $C(u,v)=u-C(u,1-v)$, for $(u,v)\in[0,1]^{2}$.
  2. (ii)
    If $(Y,Z)^{\top}$ is v-symmetric then $C(u,v)=v-C(1-u,v)$, for $(u,v)\in[0,1]^{2}$,
  3. (iii)
    If $(Y,Z)^{\top}$ is radially symmetric then $C(u,v)=u+v+C(1-u,1-v)-1$ , for $(u,v)\in[0,1]^{2}$.
Proof.
Since the copula is invariant under strictly increasing transformations of the margins, such as changes of location, we can, without loss of generality, assume that the axes and centres of symmetry are at zero and the origin. For case (i) observe that if $(Y,Z)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(Y,-Z)^{\top}$ then the joint distribution function satisfies
$F_{Y,Z}(y,z)=\mathbb{P}(Y\leqslant y,Z\leqslant z)=\mathbb{P}(Y\leqslant y,-Z\leqslant z)$ $=\mathbb{P}(Y\leqslant y)-P(Y\leqslant y,Z\leqslant-z)$
$=F_{Y}(y)-F_{Y,Z}(y,-z)$ (2)
for all $(y,z)\in\mathbb{R}^{2}$ and application of Sklar’s Theorem (Sklar,, 1959) implies that the (unique) copula $C$ must satisfy
$C(F_{Y}(y),F_{Z}(z))=F_{Y}(y)-C(F_{Y}(y),F_{Z}(-z))=F_{Y}(y)-C(F_{Y}(y),1-F_{Z}(z))$
where $F_{Y}$ and $F_{Z}$ denote the marginal distribution functions and the final equality follows from the symmetry of the marginal distribution of $Z$. Hence we obtain that $C(u,v)=u-C(u,1-v)$ at all points in the unit square. The proof of (ii) is obviously identical to (i) while (iii) is a well-known result found in Nelsen, (2006) (Theorem 2.7.3). ∎
Definition 2.
A copula $C$ satisfying $C(u,v)=u-C(u,1-v)$ for all $(u,v)\in[0,1]^{2}$ will be said to have h-symmetry. Similarly, a copula satisfying $C(u,v)=v-C(1-u,v)$ will be said to have v-symmetry. A copula satisfying $C(u,v)=u+v+C(1-u,1-v)-1$ will be said to have radial symmetry. A copula with all three forms of symmetry will be said to be jointly symmetric.
We note that joint symmetry of the copula is the bivariate special case of a property that Durante and Fuchs, (2019) call reflection invariance.
2.2 Correlation and rank correlation
Nelsen, (2006) notes that jointly symmetric distributions have zero correlation (when finite second moments permit the measure to be defined). Perhaps surprisingly, h-symmetry on its own is a sufficient condition for an absence of both correlation and rank correlation. The same is true of v-symmetry.
Proposition 1.
Proof.
We give the proof in the case of h-symmetry, the method for v-symmetry being identical. If $(Y,Z)^{\top}$ is h-symmetric about $b$ then $\mathbb{E}(Z)=b$ and we can write $\operatorname{cov}(Y,Z)=\mathbb{E}(YZ)-\mathbb{E}(Y)\mathbb{E}(Z)=\mathbb{E}(Y(Z-b))+\mathbb{E}(Yb)-\mathbb{E}(Y)b=\mathbb{E}(Y(Z-b))$. However, h-symmetry implies $\mathbb{E}(Y(Z-b))=\mathbb{E}(Y(b-Z))=-\mathbb{E}(Y,(Z-b))$ implying that $\mathbb{E}(Y(Z-b))=0$ from which the result follows. ∎
Proposition 2.
Proof.
Again it suffices to give the proof in the case of h-symmetry. Let the unique copula of $(Y,Z)^{\top}$ be denoted $C$. For Spearman’s rank correlation we need to calculate $\rho_{S}(Y,Z)=12\int_{0}^{1}\int_{0}^{1}(C(u,v)-uv)\mathrm{d}u\mathrm{d}v$. From the identity $C(u,v)=u-C(u,1-v)$ we infer that
where we have made the change of variable $y=1-v$. We must have $\int_{0}^{1}\int_{0}^{1}C(u,v)\mathrm{d}u\mathrm{d}v=\tfrac{1}{4}$ and then simple calculations give $\rho_{S}(Y,Z)=0$.
In the case of Kendall’s rank correlation we need to compute $\tau(Y,Z)=4\mathbb{E}(C(U,V))-1$ where $(U,V)^{\top}$ is a pair of variables with distribution function $C$. The identity $C(u,v)=u-C(u,1-v)$ implies that $(U,V)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(U,1-V)^{\top}$; indeed $(U,V-0.5)^{\top}\stackrel{{\scriptstyle\text{\tiny d}}}{{=}}(U,0.5-V)^{\top}$ since $(U,V)^{\top}$ is h-symmetric about 0.5. Hence we infer that
which implies that $\mathbb{E}(C(U,V))=\tfrac{1}{4}$ and $\tau(Y,Z)=0$. ∎
Despite the absence of correlation in distributions with these forms of symmetry, we will see in this paper that they can have quite strong forms of dependence.
2.3 Copula densities
We will analyse GARCH-type models in terms of their joint densities. Henceforth we assume that the generic random vector $(Y,Z)^{\top}$ has joint density $f_{Y,Z}$. It is easily inferred from Definition 1, part (i), that h-symmetry is equivalent to $f_{Y,Z}(y,b+z)=f_{Y,Z}(y,b-z)$ for all $(y,z)\in\mathbb{R}^{2}$ and some $b\in\mathbb{R}$. Similarly, v-symmetry and radial symmetry are equivalent to the density satisfying $f_{Y,Z}(a+y,z)=f_{Y,Z}(a-y,z)$ for some $a\in\mathbb{R}$ and $f_{Y,Z}(a+y,b+z)=f_{Y,Z}(a-y,b-z)$ for some $a,b\in\mathbb{R}$, respectively.
The density $c$ of the copula $C$ of $(Y,Z)^{\top}$ is given by
$c(u,v)=\frac{f_{Y,Z}\left(F_{Y}^{-1}(u),F_{Z}^{-1}(v)\right)}{f_{Y}\left(F_{Y}^{-1}(u)\right)f_{Z}\left(F_{Z}^{-1}(v)\right)}$ (3)
where $f_{Y}$ and $f_{Z}$ denote the marginal densities. The identity $C(u,v)=u-C(u,1-v)$ for h-symmetry of the copula $C$ in Definition 2 is equivalent to the identity $c(u,v)=c(u,1-v)$ for densities (symmetry in the line $v=0.5$). Similarly, v-symmetry is equivalent to $c(u,v)=c(1-u,v)$ (symmetry in the line $u=0.5$) and radial symmetry is equivalent to $c(u,v)=c(1-u,1-v)$.
The following result provides the first connection between copula densities and v-transforms. The function $\mathcal{V}(u)=|2u-1|$ appearing in this result is the first and simplest example of a v-transform, as we will discuss in Section 4.
Proposition 3.
Let $(Y,Z)^{\top}$ be a random vector with joint density $f_{Y,Z}$ and copula density $c$. Then, for all $(u,v)\in[0,1]^{2}$,
  1. (i)
    if $(Y,Z)^{\top}$ is h-symmetric about $b$, $c(u,v)=c^{*}(u,|2v-1|)$ where $c^{*}$ is the density of the copula of $(Y,|Z-b|)^{\top}$;
  2. (ii)
    if $(Y,Z)^{\top}$ is v-symmetric about $a$, $c(u,v)=c^{*}(|2u-1|,v)$ where $c^{*}$ is the density of the copula of $(|Y-a|,Z)^{\top}$.
Proof.
It suffices to show part (i). We set $\tilde{Z}=Z-b$ and observe that $(Y,\tilde{Z})^{\top}$ is h-symmetric about 0. We then find that
$F_{Y,|\tilde{Z}|}(y,z)=\mathbb{P}(Y\leqslant y,|\tilde{Z}|\leqslant z)=\mathbb{P}(Y\leqslant y,\tilde{Z}\leqslant z)-\mathbb{P}(Y\leqslant y,\tilde{Z}\leqslant-z)=F_{Y,\tilde{Z}}(y,z)-F_{Y,\tilde{Z}}(y,-z)$
and hence $f_{Y,|\tilde{Z}|}(y,z)=f_{Y,\tilde{Z}}(y,z)+f_{Y,\tilde{Z}}(y,-z)=2f_{Y,\tilde{Z}}(y,z)$ by h-symmetry about zero. Similarly, $f_{|\tilde{Z}|}(z)=2f_{\tilde{Z}}(z)$. From (3) we have
$c^{*}(u,|2v-1|)=\frac{f_{Y,|\tilde{Z}|}\left(F_{Y}^{-1}(u),F_{|\tilde{Z}|}^{-1}(|2v-1|)\right)}{f_{Y}\left(F_{Y}^{-1}(u)\right)f_{|\tilde{Z}|}\left(F_{|\tilde{Z}|}^{-1}(|2v-1|)\right)}=\frac{f_{Y,\tilde{Z}}\left(F_{Y}^{-1}(u),F_{|\tilde{Z}|}^{-1}(|2v-1|)\right)}{f_{Y}\left(F_{Y}^{-1}(u)\right)f_{\tilde{Z}}\left(F_{|\tilde{Z}|}^{-1}(|2v-1|)\right)}.$
Since $F_{|\tilde{Z}|}(z)=2F_{\tilde{Z}}(z)-1$ for $z\geqslant 0$, it follows that $F_{|\tilde{Z}|}^{-1}(q)=F_{\tilde{Z}}^{-1}((1+q)/2)$ for $q\in[0,1]$. Therefore, for $v\leqslant 0.5$, we have
$F_{|\tilde{Z}|}^{-1}(|2v-1|)=F_{|\tilde{Z}|}^{-1}(1-2v)=F_{\tilde{Z}}^{-1}(1-v)=-F_{\tilde{Z}}^{-1}(v)$
while, for $v\geqslant 0.5$, we have $F_{|\tilde{Z}|}^{-1}(|2v-1|)=F_{|\tilde{Z}|}^{-1}(2v-1)=F_{\tilde{Z}}^{-1}(v)$. Using the h-symmetry of $f_{Y,\tilde{Z}}$ and the symmetry of $f_{\tilde{Z}}$ we infer that $c^{*}(u,|2v-1|)=\tilde{c}(u,v)$ where $\tilde{c}$ is the density of the copula of $(Y,\tilde{Z})^{\top}$. The identity $c^{*}(u,|2v-1|)=c(u,v)$ follows from the fact that $(Y,Z)^{\top}$ and $(Y,\tilde{Z})^{\top}$ have the same copula. ∎
Clearly, by putting both parts of Proposition 3 together, we see that the density of the copula of a random vector $(Y,Z)^{\top}$ that is jointly symmetric about $(a,b)$ satisfies $c(u,v)=c^{*}(|2u-1|,|2v-1|)$ where $c^{*}$ is the density of the copula of $(|Y-a|,|Z-b|)^{\top}$.
3 ARCH and GARCH copulas

### v2

# GARCH copulas, v-transforms and D-vines for stochastic volatility

@@ Abstract
The bivariate copulas that describe the dependencies and partial dependencies of lagged variables in strictly stationary, first-order GARCH-type processes are investigated. It is shown that the copulas of symmetric GARCH processes are jointly symmetric but non-exchangeable, while the copulas of processes with symmetric innovation distributions and asymmetric leverage effects have weaker h-symmetry; copulas with asymmetric innovation distributions have neither form of symmetry. Since the true bivariate copulas are typically inaccessible, due to the unknown functional forms of the marginal distributions of GARCH processes, a new class of approximating copulas is proposed. These rely on copula density constructions that combine standard bivariate copula densities for positive dependence with two uniformity-preserving transformations known as v-transforms. The construction is shown to be particularly effective when applied to the density of the copula of the absolute values of a spherical t distribution. Tractable simplified D-vines incorporating the new pair copulas are developed for applications to time series showing stochastic volatility. The resulting models are shown to provide better fits to simulated data from GARCH processes, and to a dataset of financial exchange-rate returns, than have previously been obtained using vine copulas.
@@ 1 Introduction
This paper addresses the construction of copula-based models for time series displaying stochastic volatility, such as financial asset-price return series. The philosophy of the copula-modelling approach is that, by separating the issues of modelling marginal behaviour and dependence structure, we may obtain composite models that provide superior fit and predictive performance in practice. A number of authors (Loaiza-Maya et al.,, 2018; Bladt and McNeil, 2022b, ; Zhao et al.,, 2022) have suggested that simplified D-vines, a class of models based on pair or bivariate copulas, offer a promising framework for modelling volatile return series.
Finding suitable pair copulas to describe volatile return series is not straightforward, since most existing bivariate copulas are for modelling monotonic dependence relationships. The serial dependencies in return series tend to be monotonic in absolute values, but non-monotonic in the raw returns. In this paper we attempt to learn from the success of GARCH-type models in practical applications by studying their implied pair copulas. Primarily, this is a paper about the dependence structure of GARCH processes rather than empirical data, although we provide an example to show that the insights we gain can lead to highly effective D-vine models for real data.
@@
The ARCH (Engle,, 1982) and GARCH (Bollerslev,, 1986) prototypes have inspired a vast class of related GARCH-type models, including the GJR-GARCH model, which incorporates leverage effects to reflect different contributions to volatility by returns of differing signs (Glosten et al.,, 1993), power-GARCH models (Ding et al.,, 1993) and exponential GARCH models (Nelson,, 1991), to name but a few of the more important extensions. Despite their success in modelling financial data, GARCH-type models are something of a black box. The stationary distribution is not generally known in closed form, although many models give rise to distributions with Pareto tails (Mikosch and Stărică,, 2000). In particular, the serial dependence structure is not well understood. If $(X_{t})_{t\in\mathbb{Z}}$ follows a stationary GARCH-type model then the distribution of $(X_{1},\ldots,X_{n})^{\top}$ for $n\geqslant 2$ must have a copula, but this copula has no simple, accessible closed form.
@@
Vine copulas also have a large literature, including Joe, (1996), Bedford and Cooke, (2002), Kurowicka and Cooke, (2006), Aas et al., (2009) and Smith et al., (2010). The most suitable vine structure for a time series is the D-vine, which is able to describe strict stationarity of a random vector under additional translation-invariance restrictions on the vine structure (Nagler et al.,, 2022). Markov copula models (Darsow et al.,, 1992; Chen and Fan,, 2006; Beare,, 2010; Domma et al.,, 2009) are examples of stationary first-order D-vines and a number of authors have written on their higher-order and multivariate extensions (Ibragimov,, 2009; Beare and Seo,, 2015; Brechmann and Czado,, 2015; Loaiza-Maya et al.,, 2018; Nagler et al.,, 2022; Zhao et al.,, 2022).
Our goals in the present paper are threefold. First, we want to analyse the structure of the implied pair copulas in first-order GARCH-type models and to investigate how different model features affect the symmetries of these pair copulas. Second, since the true GARCH pair copulas are inaccessible for practical purposes, we want to find tractable copulas that can approximate them. Finally, with the help of the analyses in the paper, we aim to propose a class of D-vines that behave like GARCH-type models and give a first glimpse of its suitability for real data applications.
Our approach is similar to the one followed in the insightful paper of Loaiza-Maya et al., (2018), in which the authors use simulated data to reveal the cross-shaped dependence structure of the copula that describes the serial dependence structure of a classical symmetric ARCH(1) process [^1] Loaiza-Maya et al., (2018) and Smith and Maneesoonthorn, (2018) also analyse the first-order copulas of discrete-time stochastic volatility models, an alternative model class for financial asset-price returns we do not consider; they find similar cross-shaped behaviour.. They use this analysis to propose D-vines of low Markov order (1 or 5) in which the pair copulas are mixtures of positive dependence and negative dependence copulas; they show that these models give better forecasts of the conditional quantiles (value-at-risk) of exchange-rate returns than classical symmetric models of first-order GARCH type. Another paper related to our own is by Zhao et al., (2022), who argue that t copulas in D-vines can approximate the dependence structure of GARCH(1,1) and GJR-GARCH(1,1) models. Our paper also extends ideas in Bladt and McNeil, 2022b by proposing a class of models that partly overlaps with the vt-D-vines in that paper but offers additional flexibility.
@@
To explain our contributions, we note that the joint density of the distribution of $(X_{1},\ldots,X_{n})^{\top}$ in a strictly stationary process $(X_{t})_{t\in\mathbb{Z}}$, such as a stationary model of GARCH type, can be decomposed as a D-vine of the form
$f_{X_{1},\ldots,X_{n}}(x_{1},\ldots,x_{n})=\prod_{i=1}^{n}f_{X}(x_{i})\\ \prod_{k=1}^{n-1}\prod_{j=k+1}^{n}c_{k}(F_{X_{1}|\bm{X}_{[2:k]}}(x_{j-k}\mid\bm{x}_{[j-k+1:j-1]}),F_{X_{k+1}|\bm{X}_{[2:k]}}(x_{j}\mid\bm{x}_{[j-k+1:j-1]})\mid\bm{x}_{[j-k+1:j-1]})$ (1)
@@
where we use the compact notation $\bm{x}_{[l:m]}:=(x_{l},\ldots,x_{m})^{\top}$, $f_{X}$ denotes the marginal density, $F_{X_{i}|\bm{X}_{[2:k]}}$ is the conditional distribution function of $X_{i}$ given $X_{2},\ldots,X_{k}$ and $c_{k}(u,v\mid\bm{x}_{[j-k+1:j-1]})$ is the density of the copula of the conditional distribution of $(X_{1},X_{k+1})^{\top}$ given $X_{2}=x_{j-k+1},\ldots,X_{k}=x_{j-1}$. In the case $k=1$ the conditional distribution is simply the marginal distribution $F_{X}$ and $c_{1}$ is simply the density of the copula of the unconditional distribution of $(X_{1},X_{2})^{\top}$. When the process is first-order Markov, such as a model of ARCH(1) type, all of the terms involving $c_{k}(u,v\mid\cdot)$ for $k>1$ disappear.
@@
Although the densities of GARCH-type models may be written as D-vines, they may not in general be written as so-called simplified D-vines, that is models in which the conditional copula densities $c_{k}$ do not depend explicitly on the conditioning values $\bm{x}_{[j-k+1:j-1]}$ in equation (1) for $k\geqslant 2$. When D-vines are used in practical modelling applications they are always simplified D-vines. However, as pointed out by a number of authors (Haff et al.,, 2010; Stöber et al.,, 2013; Spanhel and Kurz,, 2019; Mroz et al.,, 2021), these may not be able to capture all the features present in a given joint density, such as that of a GARCH-type model.
@@
Our first contributions relate to the analysis of the GARCH pair copula densities $c_{k}$ in (1) for first-order GARCH-type models and are described in Section 2. We provide theory to show that (i) symmetric models give rise to jointly symmetric copulas, (ii) models with symmetric innovation distributions and leverage have copulas with weaker symmetry, which we call h-symmetry, and (iii) models with asymmetric innovations lose all symmetry. We confirm the theoretical insights by using numerical approximations to provide accurate contour plots of the copula densities which reveal hitherto overlooked features, such as the non-exchangeability of copulas even in the most symmetric models. The theory complements and extends Loaiza-Maya et al., (2018), who consider the copula density $c_{1}$ in the ARCH(1) model and use a simulation approach and adaptive kernel density estimation to elicit its cross-shaped form.
@@
The second body of material, found in Section 3 relates to the search for copulas that can approximate the GARCH pair copulas. We show how uniformity-preserving transformations known as v-transforms (McNeil,, 2021) play an implicit role in the copulas of classic symmetric models and in models with certain forms of asymmetry. We investigate methods of constructing copulas to approximate GARCH pair copulas using a technique we call stochastic inversion of v-transforms and we identify the copulas that provide the best fit to simulated data from GARCH-type models. We find that inverse-v-transformed versions of positive dependence copulas with upper tail dependence, such as survival Clayton, Joe, or a new copula we refer to as absolute spherical t, typically fit better than mixture copulas.
Finally, in Section 4, we propose a class of simplified D-vines that behave like GARCH models and that could be used to model real data. To support the use of simplified D-vines, we provide some evidence in Section 2.5 that the dependence of higher order conditional copulas on conditioning variables may be weak. Our simplified D-vines combine the inverse-v-transformed pair copulas developed in Section 3 with a parsimonious method of parameterization via the partial autocorrelation function of ARMA processes as proposed by Bladt et al., (2025). The resulting models generalize the vt-D-vine models of Bladt and McNeil, 2022b and provide a better fit to simulated GARCH data than the simplified D-vines of order 1 and 2 based on t copulas used in Zhao et al., (2022). A first application to a real-world dataset of daily exchange-rate returns provided and analysed by Loaiza-Maya et al., (2018) suggests that the much high Markov order of our models, and the copula constructions they contain, can lead to better-fitting models than have previously been obtained using D-vines.
2 An analysis of GARCH pair copulas